1. 42 is 70% of what number?

2. what percent of 140 is 210?
3. 6% of what number is $24?
4. what percent of 3,100 is 930?

Answer by stanbon(75887)


1. 42 is 70% of what number?
Let the number be "x"::
Equation:
42 = 0.7*x
x = 42/0.7 = 4200/7 = 600
------------------------------------
2. what percent of 140 is 210?
x/100 = 210/140
x/100 = 21/14
x/100 = 3/2 = 1.5
x = 150%
-----------------------------------
3. 6% of what number is $24?
0.06x = 24
x = 24/0.06
x = 2400/6
x = 400%
----------------------
4. what percent of 3,100 is 930?
x/100 = 930/3100
x/100 = 93/310
x/100 = 3/10 = 0.3
x = 30%
-------------------
Cheers,
Stan H.
-------------